Can you exlain the steps to find the answer of:
64=x^2+12x

Answer by algebrahouse.com(1659)

64 = x² + 12x
x² + 12x - 64 = 0 {subtracted 64 from both sides}
(x + 16)(x - 4) = 0 {factored into two binomials}
x + 16 = 0 or x - 4 = 0 {set each factor equal to 0}
x = -16 or x = 4 {solved each equation for x}

For more help from me, visit: www.algebrahouse.com